Solve the following equation for X:
logx + log(X-1)= Log(4x)

Answer options
a)X=0,5
B)X=1/4log(x-1)
C)X=5
D)there are no solutions

Answer by stanbon(75887)

Solve the following equation for X:
logx + log(X-1)= Log(4x)
========
log[x(x-1)] = log(4x)
----
x^2-x = 4x
-----
x^2-5x = 0
Factor:
x(x-5) = 0
x = 0 (extraneous)
x = 5 (only answer)
===================
Cheers,
Stan H.
===================
Answer options
a)X=0,5
B)X=1/4log(x-1)
C)X=5
D)there are no solutions